state whether each of the following is true or false for all real numbers x
a)
sq rt x^2 = x
b)
sq rt x^2 + 9 = x+3
c)
2x+y/4=x+y/2
d)
sq rt 1-x= -sq rt x-1

Answer by rajagopalan(148)

a)sq rt x^2 = x .. True. you have another answer also in -x.
b)sq rt x^2 + 9 = x+3.. False ..when RHs is squared we get x^2+9+2x ..OK
c)2x+y/4=x+y/2 True ..both sides are didvided by 2
d)sq rt 1-x= -sq rt x-1 False
since
sqrt(1-x)=sqrt(-1)*(x-1)=-i*sqrt(x-1)an imaginary number